The difference of the areas of two squares is 288 square feet. Each side of the larger square is three times the length of a side of the smaller square. Find the length of a side of each square.

Answer by mananth(16949)

smaller square side = x
larger square side = 3x
..
smaller square area = x^2
larger square area = 9x^2
difference = 9x^-x^2=8x^2
...
8x^2= 288
/8
x^2=36
x=6 feet side of smaller square
larger square side = 3x = 18feet